### v1

# Improving Retrieval-Augmented Generation through Multi-Agent Reinforcement Learning

Abstract.
Retrieval-augmented generation (RAG) is extensively utilized to incorporate external, current knowledge into large language models, thereby minimizing hallucinations. A standard RAG pipeline may comprise several components, such as query rewriting, document retrieval, document filtering, and answer generation. However, these components are typically optimized separately through supervised fine-tuning, which can lead to misalignments between the objectives of individual modules and the overarching aim of generating accurate answers in question-answering (QA) tasks. Although recent efforts have explored reinforcement learning (RL) to optimize specific RAG components, these approaches often focus on overly simplistic pipelines with only two components or do not adequately address the complex interdependencies and collaborative interactions among the modules. To overcome these challenges, we propose treating the RAG pipeline as a multi-agent cooperative task, with each component regarded as an RL agent. Specifically, we present MMOA-RAG, a Multi-Module joint Optimization Algorithm for RAG, which employs multi-agent reinforcement learning to harmonize all agents’ goals towards a unified reward, such as the F1 score of the final answer. Experiments conducted on various QA datasets demonstrate that MMOA-RAG improves the overall pipeline performance and outperforms existing baselines. Furthermore, comprehensive ablation studies validate the contributions of individual components and the adaptability of MMOA-RAG across different RAG components and datasets [^1] The code of MMOA-RAG is on https://github.com/chenyiqun/MMOA-RAG..
Retrieval-Augmented Generation; Multi-Agent Cooperation; Multi-Agent Reinforcement Learning; Multi-Module Joint Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Information systems Question answering
1. Introduction
Large Language Models (LLMs) have been widely applied to tasks such as question answering (Asai et al., 2023; Khattab et al., 2022), information retrieval (Sun et al., 2023; Chen et al., 2024), various forms of reasoning (Huang and Chang, 2022; Hao et al., 2023), and evaluation (Gong and Mao, 2023; Fu et al., 2023). However, since LLMs cannot promptly update their internal knowledge after pre-training, they remain prone to generating outdated or fabricated responses (Zhao et al., 2023). To address these challenges, Retrieval-Augmented Generation (RAG) enhances the generative capabilities of LLMs by retrieving relevant information from external knowledge sources. Recent RAG systems are often built as complex pipelines comprising multiple interconnected modules  (Gao et al., 2023), including query rewriting (Ma et al., 2023; Jiang et al., 2024), first-stage retrieval (Lewis et al., 2020; Shao et al., 2023), re-ranking (Salemi and Zamani, 2024a, b), document pre-processing (Ke et al., 2024; Li et al., 2024b), and answer generation (Shao et al., 2023; Su et al., 2024).
The complexity of RAG systems makes their optimization particularly challenging. Standard supervised fine-tuning (SFT) optimizes each module independently using human-annotated data. However, this often results in misalignment between the objectives of individual components and the system’s overarching goal of generating high-quality outputs. For example, retrieval modules are frequently trained on human-labeled relevance data to optimize metrics such as nDCG(Järvelin and Kekäläinen, 2002), but this process fails to address the disconnect between document relevance and response quality—documents with high relevance scores do not always contribute to generating accurate answers  (Cuconasu et al., 2024).
To address this issue, existing work on end-to-end optimization for RAG, such as  (Lee et al., 2019; Guu et al., 2020; Lewis et al., 2020; Salemi and Zamani, 2024b) aims to propagate rewards from the final output to intermediate modules using techniques like attention distributions (Izacard and Grave, 2020), generation probability (Lewis et al., 2020; Zamani and Bendersky, 2024), and expectation-maximization (EM) iterations (Singh et al., 2021; Salemi and Zamani, 2024a). However, earlier approaches primarily focus on simplified pipelines with only two components-a retriever and a generator-and fail to provide a generalizable framework for jointly optimizing complex systems with multiple components and richer interdependencies. More recent methods attempt to eliminate the need for module-specific rewards by leveraging algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Proximal Policy Optimization (PPO)(Schulman et al., 2017). Nonetheless, these methods still concentrate on optimizing individual RAG modules in isolation, without adequately modeling the collaborative dynamics between interacting components (Li et al., 2024b; Ma et al., 2023; Ke et al., 2024). Effectively capturing interdependencies among multiple modules and jointly optimizing complex RAG architectures remains an open research challenge.
Figure 1. The overall framework of MMOA-RAG.
In this paper, we propose a novel approach called the Multi-Module joint Optimization Algorithm (MMOA-RAG) to enable joint optimization across multiple modules in a RAG system. Our framework treats each intermediate component in the RAG pipeline as an agent and formulates the optimization process as a multi-agent collaborative reinforcement learning (RL) problem, where the agents (i.e., modules) work together to maximize a shared reward for the final outcome.
While MMOA-RAG is flexible with the choices reward function and pipeline design, in this study, we define the final reward as the correctness of the generated response, measured by the F1 score against the ground-truth answer, and apply MMOA-RAG to a RAG pipeline comprising four modules: a query rewriter, a fixed document retriever, a document selector, and an answer generator. To achieve this, we employ the Multi-Agent PPO (Yu et al., 2022) algorithm, enabling collaborative optimization within a fully cooperative setting. In this setting, the optimization objectives of all modules are aligned toward the ultimate goal of generating high-quality answers.
Compared to previous approaches for RAG, MMOA-RAG allows for the end-to-end optimization of complex RAG systems, ensuring that each module’s objectives align with the overarching goal of producing accurate responses. Additionally, unlike recent methods based on DPO (Zhu et al., 2024a, b) or PPO (Ma et al., 2023; Ke et al., 2024), our approach is more flexible for various pipeline designs and excels at fostering collaboration among multiple modules.
To demonstrate the effectiveness of the MMOA-RAG modeling and optimization approach, we conducted experiments on three publicly available QA datasets, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020) and AmbigQA (Min et al., 2020), based on Llama-3-8B-Instruct (Dubey et al., 2024). The experimental results indicate that MMOA-RAG achieves better performance than a series of existing optimization methods for RAG. Additionally, we performed extensive ablation studies to investigate the effectiveness and advantages of jointly optimizing mulitple modules in the RAG system and the generalizability of MMOA-RAG across different RAG pipelines.
Our main contributions are as follows:
  • We innovatively model RAG as a multi-agent collaborative task, treating multiple modules within the RAG pipeline as individual agents.
  • We employ a multi-agent cooperative reinforcement learning algorithm to jointly optimize a sophisticated RAG system with three modules: a query writer, a document selector, and an answer generation.
  • We conduct extensive experiments to verify and demonstrate the effectiveness of the proposed framework.
2. Related Works
2.1. End-to-end optimization in OpenQA
ORQA (Lee et al., 2019) is an open-domain QA system that learns end-to-end evidence retrieval and answer generation using only question-answer pairs, enabled by pretraining with an Inverse Cloze Task. REALM (Guu et al., 2020) is an end-to-end optimizing framework that enhances language model pre-training with a retrieval-augmented approach. Lewis et al. (2020) introduces Retrieval-Augmented Generation as RAG, a model that combines pre-trained language models with non-parametric memory for improved performance on knowledge-intensive NLP tasks. Izacard and Grave (2020) propose a knowledge distillation method to train retriever models using synthetic labels derived from reader model attention scores. Stochastic RAG (Zamani and Bendersky, 2024) introduces a novel end-to-end optimization framework for RAG through expected utility maximization.
2.2. RAG without parameters update
These methods typically involve designing a novel RAG mechanism to enhance the performance of LLMs on question answering tasks. For example, DSP (Khattab et al., 2022) leverages sophisticated interactions between retrieval and language models to address knowledge-intensive NLP tasks. FLARE (Jiang et al., 2023), an active retrieval augmented generation method, enhances text generation by dynamically retrieving relevant information throughout the process, showing superior performance across various long-form knowledge-intensive tasks. ITER-RETGEN (Shao et al., 2023) is an iterative retrieval-generation synergy method that enhances retrieval-augmented large language models by synergistically combining retrieval and generation in an iterative manner. Search-in-the-Chain (Xu et al., 2024a) is a framework that interactively enhances Large Language Models with search capabilities to improve performance on complex, knowledge-intensive tasks. SELF-RAG (Asai et al., 2023) enhances language model quality and factuality through self-reflective retrieval and generation. DRAGIN (Su et al., 2024) is a dynamic RAG framework that addresses the real-time information needs of LLMs during text generation, enhancing performance on knowledge-intensive tasks. GenGround (Shi et al., 2024) synergizes large language model knowledge with external documents to enhance multi-hop question answering through an iterative process of generating answers and grounding them in evidence. Astute RAG (Wang et al., 2024) is a approach that enhances the robustness of Retrieval-Augmented Generation for Large Language Models by adaptively integrating internal and external knowledge while resolving knowledge conflicts.
2.3. RAG with parameters update
2.3.1. Optimizing RAG with SFT
INFO-RAG (Xu et al., 2024b), an unsupervised training method, enhances the capacity of large language models to integrate and refine information from retrieved texts. LongRAG (Zhao et al., 2024b) introduces a dual-perspective retrieval-augmented generation system to enhance understanding of complex long-context knowledge for improved performance in long-context question answering tasks. In INSTRUCTRAG (Wei et al., [n. d.]), generation accuracy and trustworthiness are enhanced by explicitly denoising retrieved information through self-synthesized rationales, outperforming standard RAG approaches without additional supervision.
2.3.2. Optimizing RAG with RL
Some existing works use PPO (Schulman et al., 2017) algorithm to fine-tune LLMs. In Rewrite-Retrieve-Read framework (Ma et al., 2023), a small language model is trained with reinforcement learning to rewrite queries for RAG. BGM (Ke et al., 2024) proposes a novel bridge mechanism between retrieval model and LLMs and uses PPO to optimize the parameters of the bridge to filter for more helpful documents. SMARTRAG (Gao et al., 2024) optimizes an iterative RAG framework with reward, which includes a decision maker and a policy network. RAG-Star (Jiang et al., 2024) is a reasoning approach that combines Monte Carlo Tree Search (MCTS) to improve the complex reasoning abilities of LLMs.
Some other works use DPO (Rafailov et al., 2024) or similar alignment algorithms to optimize LLMs. A noise-filtering method (Zhu et al., 2024a) is proposed by optimizing mutual information between compressed data and output while minimizing it with the retrieved passage. ATM (Zhu et al., 2024b), an Adversarial Tuning Multi-agent system, enhances the robustness and performance of retrieval-augmented generators in question answering by iteratively tuning against an adversarial attacker agent to better discriminate useful documents and resist fabricated content. SEER (Zhao et al., 2024a) proposes a novel self-aligned evidence extraction learning framework aimed at enhancing RAG performance by optimizing the extraction of high-quality, concise, and relevant evidence. RAG-DDR (Li et al., 2024b) optimizes RAG systems by aligning data preferences between modules through DDR, resulting in enhanced performance on knowledge-intensive tasks.
3. Preliminary
3.1. Modeling RAG as Co-MARL
In this work, we conceptualize the RAG procedure within a cooperative multi-agent reinforcement learning (Co-MARL) framework. Within this framework, each module of the RAG pipeline functions as an individual RL agent. The overarching objective of this multi-agent system is to produce high-quality answers, which aligns with the individual goals of each module.
In this context, we define the tuple $\left\langle\mathcal{G},\mathcal{O},\mathcal{A},\mathcal{R}\right\rangle$, where $\mathcal{G}$ denotes the set of agents in the Co-MARL system, $\mathcal{O}$ represents the observation information available to each agent, $\mathcal{A}$ constitutes the action space accessible to each agent, and $\mathcal{R}$ is the reward shared among all agents. The ultimate aim is to maximize this shared reward, thereby achieving higher evaluation metrics and enhancing the overall performance of the RAG system.
3.2. MAPPO Algorithm
In this paper we utilizes Multi-Agent PPO (MAPPO) (Yu et al., 2022), which is an extension of the PPO algorithm (Schulman et al., 2017) for multi-agent environments, to optimize the policy for each agent in the Co-MARL framework. In fully cooperative settings, unlike PPO, which focuses on single-agent scenarios with individual reward, MAPPO employs a shared global reward to promote cooperation among all agents. Additionally, a further distinction between MAPPO and PPO lies in the input received by their critic models. In PPO, the critic model is limited to the agent’s observation information, whereas MAPPO’s global critic model can access comprehensive global information, which enables the global critic model to more accurately estimate the state-value function.
4. Method
4.1. Overall of MMOA-RAG
RAG systems typically follow a modular architecture composed of multiple interconnected components. Figure 1 illustrates the architecture of our MMOA-RAG framework, which consists of four primary modules: the Query Rewriter, the Retriever, the Selector, and the Generator:
  • Query Rewriter reformulates the initial query $q$, which may be too complex or ambiguous to resolve with a single retrieval, into a set of sub-questions denoted as $subq$.
  • Retriever retrieves relevant documents from the corpus for each sub-questions, respectively, and outputs a set of candidate document $D$.
  • Selector further filters $D$ to obtain a subset of documents $D_{\text{selected}}$ that is useful for generating the final answer to the initial query $q$.
  • Generator leverages $D_{\text{selected}}$ to generate the predicted answer $Ans_{\text{predict}}$ to the initial question.
Since the Query Rewriter, Selector, and Generator modules can all be implemented using LLMs, they can be treated as RL agents (Ouyang et al., 2022), enabling parameter updates through reward signals. To optimize computational efficiency, these three modules can share the same LLM. Additionally, given the difficulty of modeling the Retriever module as an RL agent, we consider the Retriever to be part of the environment  [^2] Recent studies in generative IR (see  Li et al. (2024a) for a survey) have explored using generative models for retrieval. But we choose a more traditional dense retrieval model (Izacard et al., 2021) as the first-stage retriever and leave the optimization of the first-stage retriever for future work..
The focus of the MMOA-RAG framework is on the collaborative optimization of multiple modules to align their individual optimization objectives with the ultimate goal of generating high-quality answers. We can use metrics derived from the Generator’s predicted answer $Ans_{\text{predict}}$, such as the F1 score, as a shared reward $R_{\text{shared}}$. Given the fully cooperative nature of the modules in the RAG system, the shared reward $R_{\text{shared}}$ can be used to train all agents, a common approach in existing MARL literature (Rashid et al., 2018; Yu et al., 2022; Chen et al., 2022). Additionally, to ensure training stability and accelerate convergence in the multi-agent system, we design penalty terms $P_{QR}$, $P_{S}$, and $P_{G}$ for each agent. A more detailed explanation will be provided in Section 4.2.
4.2. Detailed Configuration for Each Agent
[…]
4.2.1. Elements of Query Rewriter
Observation of Query Rewriter is defined as Equation (1), which contains prompt of Query Rewriter $Prompt_{QR}$ and the initial question $q$.
(1) $O_{QR}=\left\{Prompt_{QR},q\right\}$
Action Space of Query Rewriter corresponds to the vocabulary of LLMs $\mathcal{V}$.
(2) $A_{QR}=\mathcal{V}$
Reward Function of the Query Rewriter is defined as shown in Equation (3). Here, $R_{\text{shared}}$ can be the metric for the final answer, depicted as the yellow section in Figure 1. In this paper, we utilize the F1 score of the predicted answer, $Ans_{\text{predict}}$, as the shared reward. The term $P_{QR}$ serves as a penalty to discourage the Query Rewriter from generating an excessive number of sub-questions during training. Specifically, $P_{QR}$ is assigned a value of -0.5 if the number of sub-questions exceeds four, and it is set to 0 if the number of sub-questions is four or fewer.
(3) $R_{QR}=R_{\text{shared}}+P_{QR}$
4.2.2. Elements of Selector
Observation of Selector is defined as Equation (4), which contains prompt of Selector $Prompt_{S}$, the initial question $q$ and the candidate documents set $D$ with $K$ documents.
(4) $O_{S}=\left\{Prompt_{S},q,D\right\}$
Action Space of Selector only comprises of several words as Equation (5). Since the function of the Selector is to output the IDs of candidate documents helpful to answering the initial question $q$, the action space is constrained to this limited set of words. This constraint can significantly reduce the exploration space of the Selector and provide a more stable training process.
(5) $A_{S}=\left\{\text{"{0}"},\text{"{1}"},\ldots,\text{"{K-1}"},\text{"{Document}"},\text{"{,}"}\right\}$ ” }
Reward Function of Selector also contains two terms, which are $R_{\text{shared}}$ and $P_{S}$. And $P_{S}$ is a penalty term designed to prevent the Selector from generating duplicate document IDs and from outputting IDs that do not conform to the specified format (e.g., Document0,Document3,Document9). When the Selector outputs duplicate document IDs or fails to adhere to the specified format, $P_{S}$ is set to -1; otherwise, $P_{S}$ is set to 0.
(6) $R_{S}=R_{\text{shared}}+P_{S}$
4.2.3. Elements of Generator
Observation of Generator is in Equation (7), which contains prompt of Generator $Prompt_{G}$, the initial question $q$ and the selected candidate documents set $D_{\text{selected}}$ given by Selector.
(7) $O_{G}=\left\{Prompt_{G},q,D_{\text{selected}}\right\}$
Action Space of Generator $A_{G}$ is the same as Query Rewriter.
(8) $A_{G}=A_{QR}=\mathcal{V}$
Reward Function of Generator contains $R_{\text{shared}}$ and penalty term $P_{G}$, which is used to constrain the model from generating excessively long content. When the generated answer exceeds a certain length, $P_{G}$ is set to -0.5; otherwise, it is set to 0.
(9) $R_{G}=R_{\text{shared}}+P_{G}$
4.3. Warm Start with SFT
In preparation for joint optimization of multiple modules using Multi-Agent PPO, it is essential to perform a warm start for each trainable module. The warm start enables the model to better adhere to instructions across diverse tasks and reduces the exploration space during MARL joint training, thereby enhancing the efficiency of exploration and exploitation. Within the MMOA-RAG framework, there are three trainable modules: the Query Rewriter, the Selector, and the Generator. Consequently, it is necessary to construct training data for the SFT of each corresponding task.
4.3.1. Query Rewriter
In Rewrite-Retrieve-Read (Ma et al., 2023), a small language model was trained using PPO to effectively rewrite queries for RAG. Building on this approach, we utilize the publicly available query rewriting data from Rewrite-Retrieve-Read as the SFT dataset to warm start the Query Rewriter in MMOA-RAG.
4.3.2. Selector
The task of the Selector is to choose a subset $D_{\text{selected}}$ that are helpful for answering a question from a given set $D$ with $K$ candidate documents. The output format of the Selector is the IDs of the documents in $D_{\text{selected}}$ (e.g., Document0, Document4, Document6, Document7), as shown in the prompt of Selector in Table 1. Therefore, to construct SFT data for the Selector, the ground truth should be the IDs of documents that are truly useful for answering the question. One method to obtain the ground truth is to employ advanced LLMs, such as GPT-4o, to provide the ground truth. However, we have found that this approach does not yield results as good as expected. Additionally, BGM (Ke et al., 2024) introduced and optimized a bridge module which is similar to the Selector module. They proposed a method called synthesis silver passage sequence (Synthesis SPS) to construct the ground truth for SFT data. However, the Synthesis SPS method requires examining each candidate document $d_{i,j}\in D_{i}$ (candidate documents of question $i$), invoking the LLM for each check, and comparing the utility values before and after the check, making it a complex and costly method.
Table 1. The prompt of Selector agent.
system: You are a helpful, respectful and honest assistant. Your task is to output the IDs of the candidate Documents (0,1,2,…,K-1) which are helpful in answering the Question. assistant: Okay, I will provide the ID of candidate Documents which are helpful in answering the Question.
user: Question: {content of Question}
Document0: {content of Document0}
Document(K-1): {content of Document(K-1)}
assistant: OK, I received the Question and the candidate Documents.
user: Now, output the IDs of the candidate Documents (0,1,2,…,K-1) which are helpful in answering the Question: {content of Question}, for example, in the following format: Document0,Document4,Document6,Document7.
We propose a convenient heuristic approach for constructing SFT data, aimed at LLMs to effectively follow instructions and output in the desired format. As illustrated in Figure 2, for a given question $q_{i}$ and its golden answer, there are $K$ candidate documents denoted as $d_{i,j}$, where $j\in\{0,1,\cdots,K-1\}$. First, by removing certain insignificant stop words and punctuation marks from $q_{i}$ and its golden answer, and converting the words to lowercase, we obtain the set $Set_{q_{i}}$. Similarly, we perform the same operation on the $K$ candidate documents $d_{i,j}$ to obtain $Set_{d_{i,j}}$. Finally, if any word from $Set_{q_{i}}$ appears in $Set_{d_{i,j}}$, the ID of corresponding document $j$ is included in the final output as the Label of SFT. With this approach, we can rapidly and cost-effectively construct the Selector’s ground truth labels during the SFT stage. Given our focus on the subsequent joint optimization of multiple modules, this straightforward data construction method can adequately meet our requirements.
Figure 2. The convenient approach to construct the SFT data for Selector.
4.3.3. Generator
The Generator is responsible for producing the final answer, $Ans_{\text{predict}}$, based on the $D_{\text{selected}}$ provided by the Selector. Therefore, the ground truth for the SFT data of Generator is the golden answer $Ans_{\text{golden}}$.
With these approaches, the SFT data used for the warm start training of these three modules—Query Rewriter, Selector, and Generator—can be obtained. All modules can be fine-tuned using the typical loss function of SFT presented in Equation (10).
(10) $\mathcal{L_{\text{SFT}}}(\theta)=-\sum_{n=1}^{N}\log P(Y_{i}\mid X_{i};\theta)$
In Equation (10), $N$ represents the number of samples in the SFT dataset, while $\theta$ denotes the parameters of the LLM. The variable $X_{i}$ corresponds to the input content of each module. Meanwhile, $Y_{i}$ signifies the output content of each module.
4.4. Multi-Agent Optimization
After undergoing SFT, the LLM demonstrates an improved ability to follow instructions while executing the functions of Query Rewriter, Selector, and Generator. The RAG system also achieves relatively satisfactory warm-start performance. To further enhance the performance of the RAG system, which is modeled as a fully cooperative multi-agent system, it is crucial to conduct joint training of multiple agents to strengthen collaboration among them.
[…]
In the multi-agent optimization process, there are three models to consider: the Actor model, the Critic model, and the SFT model. The parameters for these models are denoted as $\theta$, $\phi$, and $\theta_{\text{SFT}}$, respectively. The role of the Actor model is to provide the response $Answer_{i}$ based on the observation $O_{i}$ for each agent $i$. The Critic model is responsible for estimating the state-value function $V_{\phi}^{i,t}$, which is a classic setup in Actor-Critic architecture within RL algorithms. The SFT model serves as a baseline for the Actor model, similar to InstructGPT (Ouyang et al., 2022). The objective is to update the parameters of both the Actor and Critic models. The overall loss function, $\mathcal{L}(\theta,\phi)$, consists of two terms: $\mathcal{L}_{\text{Actor}}(\theta)$ and $\mathcal{L}_{\text{Critic}}(\phi)$:
(11) $\mathcal{L}(\theta,\phi)=\mathcal{L}_{\text{Actor}}(\theta)+\alpha*\mathcal{L}_{\text{Critic}}(\phi)$
The Actor loss function presented in Equation (12) is similar to that used in the typical single-agent PPO (Schulman et al., 2017) algorithm. The primary difference is that multiple agents are being optimized. In Equation (12), $i\in\mathcal{G}$ denotes the three agents: Query Rewriter, Selector, and Generator. The term $r_{t}^{i}$ in Equation (13) denotes the importance sampling ratio, which measures the difference between the new and old policies. The expression $\hat{A}_{\pi_{\theta}}^{i,t}$ in Equation (14) is the advantage function, estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015). The variable $\delta_{t}^{i}$ in Equation (15) is known as the temporal difference (TD) error at time step $t$.
(12) $\mathcal{L_{\text{Actor}}}(\theta)=\sum_{i}\sum_{t}\min\left(r^{i}_{t}\hat{A}_{{\pi}_{\theta}}^{i,t},\ \text{clip}\left(r^{i}_{t},1-\epsilon,1+\epsilon\right)\hat{A}_{{\pi}_{\theta}}^{i,t}\right)$
(13) $r^{i}_{t}=\frac{\pi_{\theta}(a^{i}_{t}\mid s^{i}_{t})}{\pi_{\theta_{\text{old}}}(a^{i}_{t}\mid s^{i}_{t})}$
(14) $\hat{A}_{{\pi}_{\theta}}^{i,t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{i}_{t+l}$
(15) $\delta^{i}_{t}=R(s^{i}_{t},a^{i}_{t})+\gamma V_{\phi}(s^{i}_{t+1})-V_{\phi}(s^{i}_{t})$
Similar to InstructGPT (Ouyang et al., 2022), the final reward function $R(s^{i}_{t},a^{i}_{t})$ is defined in Equation (16). The distinction is that our approach does not require a trained reward model, as we use the evaluation metric (F1 score) of the predicted answers $Ans_{\text{predict}}$ of Generator as the shared reward $R_{\text{shared}}$ for all agents. The penalty term $P_{i}$ can also be easily obtained from the output of each agent, as introduced in Section 4.2. The components $R_{i}$ in Equation (16) are defined in Equations (3), (6), or (9). And $Answer_{i}$ represents the output generated by each agent $i$ based on its individual observation $O_{i}$.
[…]
The loss function of the Critic model, as shown in Equation (17), employs a clipping operation similar to the Actor model. Here, $\Delta V_{i,t}=V_{\phi}^{i,t}-V_{\text{target}}^{i,t}$, where $V_{\phi}^{i,t}=V_{\phi}(s_{t}^{i})$. The term $V_{\text{target}}^{i,t}$ represents the cumulative return and $s_{t}^{i}$ is the state-value function.
(17) $\mathcal{L}_{\text{Critic}}(\phi)=\sum_{i}\sum_{t}\max\left[(\Delta V_{i,t})^{2},\left(\text{clip}\left(V_{\phi}^{i,t},V_{\phi_{\text{old}}}^{i,t}\pm\epsilon\right)-V_{\text{target}}^{i,t}\right)^{2}\right]$
Initialize: The parameters of the Actor model $\theta$, the Critic model $\phi$, the SFT model $\theta_{\text{SFT}}$, and a replay buffer $\mathcal{M}=\varnothing$.
Inputs: Dataset with initial questions $q$ and corresponding golden answers $Ans_{\text{golden}}$
for $epoch\leftarrow 1$ to $N\_epoch$ do
       for $batch\leftarrow 1$ to $N\_batch$ do
             // Collect Rollout
             for each question $q\in batch$ do
                   // Query Rewriter (QR)
                   Construct observation $O_{QR}$ according to Equation (1)
                   Get sub-questions $subq$ for initial question $q$
                   Calculate the penalty term of Query Rewriter $P_{QR}$
                   // Retriever
                   Retrieve $K$ candidate documents to construct $D$ // Selector (S)
                   Construct observation $O_{S}$ according to Equation (4)
                   Select $IDs$ of helpful documents, and get $D_{\text{selected}}$
                   Calculate the penalty term of Selector $P_{S}$
                   // Generator (G)
                   Construct observation $O_{G}$ according to Equation (7)
                   Predict the $Ans_{\text{predict}}$ to initial question $q$
                   Calculate the penalty term of Generator $P_{G}$
                   // Getting Reward and Storing Tuple
                   Calculate the F1 score of $Ans_{\text{predict}}$ as the shared reward $R_{\text{shared}}$
                   Get reward for each agent $R_{i},i\in\{QR,S,G\}$, according Equation (3), (6) and (9)
                   Store tuple $\mathcal{T}=\left((O_{QR},subq,R_{QR}),(O_{S},IDs,R_{S}),(O_{G},Ans_{\text{predict}},R_{G})\right)$ in the replay buffer $\mathcal{M}$
            // Policy and Value Optimization
             for each question $q\in batch$ do
                   Compute the advantage function $\hat{A}_{{\pi}_{\theta}}^{i,t}$ using GAE
                   Calculate the loss of the Actor $\mathcal{L}_{\text{Actor}}(\theta)$ and Critic model $\mathcal{L}_{\text{Critic}}(\phi)$
                   Update the parameters of models through the overall loss function $\mathcal{L}(\theta,\phi)$ in Equation (11)
            Clear the replay buffer $\mathcal{M}$ to $\varnothing$
Output : Well-trained Actor model with parameters $\theta_{\text{trained}}$
Algorithm 1 The Training Process of Multi-Agent Optimization
The pseudocode for multi-agent optimization based on MAPPO is shown in Algorithm 1, which corresponds to the overall framework of MMOA-RAG depicted in Figure 1. For a specific question, the first step is to execute the Collect Rollout process. This process involves passing through the Query Rewriter, Retriever, Selector, and Generator, and the computed tuple $\mathcal{T}=\left((O_{QR},subq,R_{QR}),(O_{S},IDs,R_{S}),(O_{G},Ans_{\text{predict}},R_{G})\right)$ is stored in the replay buffer $\mathcal{M}$. Next, the Policy and Value Optimization process is executed where the GAE is used to estimate the advantage function $\hat{A}_{{\pi}_{\theta}}^{i,t}$. Subsequently, the overall loss function $\mathcal{L}(\theta,\phi)$ is calculated, and the parameters of both the Actor and Critic models are updated. Additionally, to accelerate the entire training process, we can run a minibatch in parallel. Ultimately, we obtain a well-trained Actor model used for subsequent inference and evaluation.
5. Experiments
Our experiments mainly aim to explore the following research questions:
  • RQ1. How does our MMOA-RAG perform compared to existing RAG optimization methods?
  • RQ2. What are the results of the ablation study on MMOA-RAG, specifically, is joint optimization of multiple modules more effective?
  • RQ3. Does the joint optimization method of MMOA-RAG exhibit generalizability across different RAG systems?
  • RQ4. How does the MMOA-RAG method perform in out-of-domain scenarios, i.e., what is its generalization capability?
5.1. Experimental Settings
5.1.1. Datasets and Evaluation
We conducted experiments using MMOA-RAG alongside various baseline models across three open-domain QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and AmbigQA (Min et al., 2020). The candidate documents are all retrieved from Wikipedia passages for three datasets. We employ three key evaluation metrics—Accuracy, Exact Match (EM), and F1 score—to assess the performance of the RAG methods.
5.1.2. Baselines
The methods detailed below serve as baseline models:
LLM w/o RAG: This approach answers questions solely based on the internal knowledge embedded within LLMs, without employing any retrieval mechanisms.
Vanilla RAG w/o train: This method leverages a retrieval model to obtain relevant documents, thereby augmenting the LLM’s internal knowledge with external sources. Here, the LLM remains in a pre-trained state and has not undergone any fine-tuning.
Vanilla RAG w SFT: Building on the Vanilla RAG framework, this variant involves a fine-tuned LLM to improve the integration of retrieved external knowledge with the LLM’s internal knowledge, potentially enhancing the quality of final answers.
SELF-RAG (Asai et al., 2023): This innovative framework advances LLM performance by incorporating both adaptive retrieval mechanisms and self-reflection processes, aiming to produce precise and dependable answers.
RetRobust (Yoran et al., 2023): This approach fortifies the RAG architecture against irrelevant contexts, thereby boosting its effectiveness in open-domain question-answering scenarios.
Rewrite-Retrieve-Read (Ma et al., 2023): A small-scale query rewriter model is trained using reinforcement learning, optimizing the interaction between retrieval and answer generation.
BGM (Ke et al., 2024): Utilizing PPO, this method trains a bridge component to filter and identify documents that are more likely to be helpful, thus refining the quality of the retrieved context.
5.1.3. Implementation Details
We utilize Contriever (Izacard et al., 2021) as the Retriever. Regardless of how many sub-questions $subq$ the Query Rewriter generates from the initial question $q$, the Selector consistently receives a fixed set of $K=10$ documents as input. For example, if the Query Rewriter yields 2 sub-questions, each sub-question is used for retrieval, with the top-5 documents from each retrieval being selected as part of the candidate documents $D$ for the Selector. Furthermore, it is important to emphasize that we do not utilize any support facts or positive passages  [^3] Some QA datasets inherently include supportive texts that aid in answering questions. And some studies incorporate these supportive texts alongside retrieved candidate documents as input to LLMs for answer prediction, which significantly enhances answer quality. However, we adhere to the natural RAG process by using only the candidate documents provided by the retriever for answer prediction, as annotated supportive texts are not present in the natural RAG workflow. that come with the official datasets to generate answers. Instead, we only use the $K$ candidate documents from the retrieval model as external knowledge for answer generation.
Besides, we employ Llama-3-8B-Instruct (Dubey et al., 2024) as the foundational LLM for the baselines and MMOA-RAG. Building on the PPO code from LLama-Factory [^4] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024), we have developed MMOA-RAG, which optimizes the RAG multi-agent system using Multi-Agent PPO. And the critical hyperparameters of MMOA-RAG are detailed in Table 2.
Table 2. Key hyperparameters in the training process of MMOA-RAG.
Name Explanation Values
$\beta_{max}$ Maximum $\beta$ in Equation (16) 0.2
$\beta_{min}$ Minimum $\beta$ in Equation (16) 0.06
$\gamma$ Key hyperparameter in GAE 1.0
$\lambda$ Key hyperparameter in GAE 0.95
$\epsilon$ Clip range in MAPPO 0.2
$\alpha$ Coefficients in Equation (11) 0.1
$lr$ Maximum learning rate 2e-5
bueffer_size Buffer size in MAPPO 128
[…]
Table 3. Performance for different methods across datasets. All the results in this table are obtained using Llama-3-8B-Instruct as the backbone. In each dataset, the highest baseline value is underscored. The final row $\Delta$ displays the improvement of MMOA-RAG over the best baseline.
Methods HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
LLM w/o RAG 25.08 21.31 31.18 27.78 23.68 29.47 27.21 20.96 33.42
Vanilla RAG w/o train 27.99 20.62 30.67 31.94 13.91 22.84 31.09 22.42 33.56
Vanilla RAG w SFT 36.18 32.30 44.49 39.47 38.28 43.36 34.41 30.74 44.36
SELF-RAG (Asai et al., 2023) 30.42 27.77 38.93 36.32 35.39 38.86 28.35 25.70 39.04
RetRobust (Yoran et al., 2023) 37.69 34.60 46.49 41.02 39.73 44.51 35.13 32.37 44.78
Rewrite-Retrieve-Read (Ma et al., 2023) 38.03 33.93 46.32 40.40 39.17 44.17 35.94 31.90 45.92
BGM (Ke et al., 2024) 36.05 32.76 44.54 39.61 38.61 43.29 36.01 32.53 45.76
MMOA-RAG (ours) 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +1.12 +1.55 +1.80 +1.71 +1.79 +1.89 +2.84 +2.22 +2.67
5.2. Comparisons with Other Methods
We conducted a comparative analysis of MMOA-RAG against multiple baselines, with the results presented in Table 3. To ensure fairness in comparison, all baselines were re-implemented according to the settings delineated in Section 5.1.3. Each method utilized Llama-3-8B-Instruct as the backbone architecture. Within these methods, the untuned modules employed the pre-trained version of Llama-3-8B-Instruct, whereas the trainable modules were derived through specific SFT processes on Llama-3-8B-Instruct. Notably, in the Rewrite-Retrieve-Read framework, the query rewrite module, which is trainable, was optimized using the PPO algorithm applied to Llama-3-8B-Instruct; meanwhile, answer generation was implemented based on the SFT-refined backbone. Regarding the BGM method, we rebuilt the bridge to connect the retrieval model and the generation model, leveraging Llama-3-8B-Instruct, with this bridge being trained using the PPO algorithm. The generation model for BGM was similarly obtained from the SFT-refined backbone.
Firstly, as shown in Table 3, MMOA-RAG demonstrates superior performance across all metrics and datasets, highlighting its effectiveness. Additionally, it is noteworthy that Vanilla RAG w/o train achieves comparable results to LLM w/o RAG across various metrics. This observation suggests that the pre-trained Llama-3-8B-Instruct struggles to effectively leverage external knowledge for answer generation, likely due to the absence of RAG-related tasks in its pre-training process, which limits its external knowledge utilization. In contrast, Vanilla RAG w SFT exhibits substantial improvements over Vanilla RAG w/o train across all evaluation metrics. This indicates that the SFT-enhanced Llama-3-8B-Instruct is adept at utilizing external knowledge, successfully extracting valuable information from noisy candidate documents to enhance the quality of generated answers.
The Rewrite-Retrieve-Read and BGM approaches enhance Vanilla RAG by respectively integrating a query rewrite module and a bridge module, each of which is trained using the PPO algorithm. As indicated in Table 3, on the multi-hop datasets HotpotQA and 2WikiMultihopQA, Rewrite-Retrieve-Read surpasses BGM, suggesting that the inclusion of a query rewrite module is more effective than adding a bridge module for these multi-hop datasets. Conversely, on the single-hop dataset AmbigQA, the performance of Rewrite-Retrieve-Read and BGM is relatively similar. Our MMOA-RAG can be conceptualized as augmenting Vanilla RAG by integrating both a Query Rewriter and a Selector, whose roles are akin to the query rewrite module in Rewrite-Retrieve-Read and the bridge module in BGM. The primary advantage of MMOA-RAG lies in its simultaneous optimization of the Query Rewriter, Selector, and Generator modules. This is achieved by aligning the objectives of these modules with the goal of generating higher-quality answers via MAPPO. The experimental results presented in Table 3 further illustrate that MMOA-RAG significantly outperforms Rewrite-Retrieve-Read, BGM, and other baselines.
The results in Table 3 and the analysis in Section 5.2 jointly answer the RQ1.
5.3. Ablation Experiments on the Optimization of Different Agents
[…]
Methods HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
MMOA-RAG w/o G 38.28 35.19 46.91 41.93 40.80 45.62 37.92 33.65 47.77
MMOA-RAG w/o S 38.81 35.27 47.61 42.47 40.99 45.95 38.69 34.38 48.23
MMOA-RAG w/o QR 37.26 34.66 46.05 42.20 40.82 45.70 38.64 34.34 48.14
MMOA-RAG 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
Table 5. Generality Experiments on RAG Systems with Varying Module Configurations. In the second column, SFT and MAPPO refer to the current module configuration following the warm start training stage (Section 4.3) and the MAPPO joint optimization training stage (Section 4.4), respectively. The symbol $\Delta$ signifies the enhancement achieved in the MAPPO stage relative to the SFT stage.
Modules Training Stage & Delta HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
QR+S+G SFT 36.00 33.04 44.69 39.54 38.50 42.97 36.55 32.60 46.71
MAPPO 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +3.15 +3.11 +3.60 +3.19 +3.02 +3.43 +2.30 +2.15 +1.88
S+G SFT 34.25 32.18 43.14 38.93 37.97 42.40 35.85 32.35 45.82
MAPPO 38.23 34.85 47.07 41.79 40.57 45.25 37.60 33.90 47.19
$\Delta$ +3.98 +2.67 +3.93 +2.86 +2.60 +2.85 +1.75 +1.55 +1.37
QR+G SFT 36.76 32.78 45.00 39.15 37.89 42.91 35.50 31.50 45.31
MAPPO 38.90 35.89 47.94 42.43 41.01 46.19 37.65 33.50 47.53
$\Delta$ +2.14 +3.11 +2.94 +3.28 +3.12 +3.28 +2.15 +2.00 +2.22
As illustrated in Table 4, the complete version of MMOA-RAG, where all three modules are jointly optimized, delivers the highest performance. This underscores the effectiveness of multi-agent joint optimization within the RAG system and validates the importance of optimizing multiple modules concurrently. Additionally, the MMOA-RAG w/o S variant achieves the best performance among the three ablation configurations. The Selector’s primary function is to refine the candidate document set $D$, yielding a higher-quality subset $D_{\text{selected}}$, which enhances the Generator’s ability to produce a superior answer $Ans_{\text{predict}}$. However, through the joint optimization process with MAPPO, the Generator acquires some denoising capabilities. Consequently, satisfactory results can be achieved even when the Selector is not optimized during joint optimization.
Figure 3. Ablation experiments on AmbigQA dataset. The horizontal axis represents the number of training samples, while the vertical axis denotes the shared reward $R_{\text{shared}}$ (F1 score) during the training process.
We also present the trajectory of the shared reward $R_{\text{shared}}$ during the training process based on ablation experiments conducted on the AmbigQA dataset, as shown in Figure 3. From Figure 3, it is evident that the reward curve for MMOA-RAG demonstrates the fastest convergence rate and achieves the highest final convergence value. This underscores the effectiveness of joint optimization across multiple modules in significantly and efficiently enhancing the performance of the RAG system. Furthermore, the training curve for MMOA-RAG w/o G in Figure 3 is noticeably slower compared to other algorithms, and the test results for MMOA-RAG w/o G on the AmbigQA dataset, as shown in Table 4, are the poorest. These findings suggest that the Generator module is the most critical component for the single-hop AmbigQA dataset.
The results in Table 4 and Figure 3 answer the RQ.2 that it is more effective to optimize multiple modules in a RAG system simultaneously.
5.4. Generality Experiments on RAG Systems with Varying Module Configurations
In this section, we evaluate the performance of MMOA-RAG in optimizing RAG systems with different numbers of agents, as detailed in Table 5. In Table 5, QR+S+G represents the RAG framework depicted in Figure 1, illustrating a multi-agent system composed of three agent, Query Rewriter, Selector and Generator. The configuration S+G results from omitting the Query Rewriter agent, relying solely on the initial question $q$ for retrieval, thereby configuring the RAG system as a two-agent (Selector and Generator) system. Conversely, QR+G denotes the exclusion of the Selector agent, forming a RAG pipeline consisting of two agents, Query Rewriter and Generator. The second column of Table 5 specifies that SFT refers to the warm start of all agents in the corresponding RAG system through supervised fine-tuning, while MAPPO refers to the joint optimization of all agents built upon SFT utilizing the MAPPO framework. The notation $\Delta$ is used to denote the performance enhancement achieved by MAPPO compared to SFT.
The experimental results in Table 5 reveal that RAG systems optimized using joint MAPPO consistently outperform those using only SFT across all datasets. This finding underscores the robust generalizability of the MMOA-RAG joint optimization approach, yielding significant performance improvements across diverse RAG configurations. Notably, the performance gains from MAPPO over SFT are approximately three percentage points on multi-hop datasets such as HotpotQA and 2WikiMultihopQA, while improvements on the single-hop dataset AmbigQA are around two percentage points. This difference may stem from the greater complexity inherent to multi-hop datasets, which potentially exacerbates misalignment among different modules during the SFT stage. These results further highlight the necessity of multi-module joint optimization, especially in the context of more challenging multi-hop datasets.
The results presented in Table 5 demonstrate the effectiveness of MMOA-RAG in optimizing various RAG systems across different configurations, thereby answering RQ.3.
5.5. Out-of-Domain Experiments
We also conducted out-of-domain (OOD) experiments to evaluate the generalization capabilities of MMOA-RAG compared to the baselines. We trained LLM on HotpotQA dataset and evaluated on the AmbigQA dataset. The experimental results are shown in Table 6.
Table 6. Out-of-domain experimental results: The model is trained on the HotpotQA dataset and tested on the AmbigQA dataset.
Methods Acc EM F1
SELF-RAG (Asai et al., 2023) 26.70 24.25 36.38
RetRobust (Yoran et al., 2023) 34.19 31.75 44.08
Rewrite-Retrieve-Read (Ma et al., 2023) 33.91 30.73 43.61
BGM (Ke et al., 2024) 32.58 29.77 42.07
MMOA-RAG (ours) 35.45 32.43 45.62
From Table 6, it is evident that MMOA-RAG demonstrates superior performance in the OOD experiments, underscoring its notable generalization capabilities and effectively answering RQ.4. Additionally, it is noteworthy that RetRobust outperforms all other baselines. This can be attributed to its strategy of integrating both relevant and irrelevant data during the SFT process, which significantly enhances the robustness and generalization abilities of the RAG system.
6. Conclusions
In this paper, we model the RAG system as a multi-agent collaborative task, wherein we consider the Query Rewriter, Selector, and Generator modules as learnable RL agents. We employ a multi-agent reinforcement learning algorithm to jointly optimize these agents, aligning the optimization goals of multiple modules with the ultimate objective of generating high-quality answers.
Our experiments demonstrate the effectiveness of our modeling approach and joint optimization method. Comprehensive ablation studies confirm the necessity and generality of multi-module joint optimization, establishing MMOA-RAG as an effective approach for optimizing RAG systems.